### v1

# Field-sensitive dislocation bound states in two-dimensional $d$-wave altermagnets

Abstract
When a two-dimensional $d$-wave altermagnet is grown on a substrate, the interplay of momentum-dependent spin splittings arising from altermagnetism and Rashba spin-orbit coupling gives rise to a nodal band structure with band degeneracies enforced by a $C_{4z}\mathcal{T}$ symmetry. If we break the $C_{4z}\mathcal{T}$ symmetry by an exchange field, the band degeneracies are found to be immediately lifted, leading to a topological band structure characterized by nontrivial strong and weak topological indices. Remarkably, both the strong topological index and the $Z_{2}$-valued weak topological indices depend sensitively on the direction of the exchange field. As a consequence of the bulk-defect correspondence, we find that the unique dependence of weak topological indices on the exchange field in this system dictates that the presence or absence of topological bound states at lattice dislocations also depends sensitively on the direction of the exchange field. When the substrate is an $s$-wave superconductor, we find that a similar dependence of band topology on the exchange field gives rise to field-sensitive dislocation Majorana zero modes. As topological dislocation bound states are easily detectable by scanning tunneling microscopy, our findings unveil a promising experimental diagnosis of altermagnetic materials among an ever growing list of candidates.
I Introduction
Altermagnetism (AM) has attracted increasing interest recently as a collinear magnetic order with salient properties distinct from the conventional collinear ferromagnetism and antiferromagnetismŠmejkal et al. (2022a, b, 2020); Hayami et al. (2019, 2020); Yuan et al. (2020, 2021); Mazin et al. (2021); Shao et al. (2021); Ma et al. (2021); Liu et al. (2022); Feng et al. (2022); Gonzalez Betancourt et al. (2023); Mazin (2023); Turek (2022); Hariki et al. (2023); Bhowal and Spaldin (2024); Zhou et al. (2024); Han et al. (2024); Chen et al. (2023); Xiao et al. (2023); Jiang et al. (2023). In real space, the magnetic moments of an altermagnet are collinearly arranged to form a Néel order, just like an antiferromagnet. However, unlike the antiferromagnets the two sublattices with opposite magnetic moments in a magnetic unit cell of the altermagnet cannot be mapped to each other by the combined symmetry operations of time reversal and inversion/translation. Instead, they are mapped to each other by the combined symmetry operations of time reversal and rotation/mirrorŠmejkal et al. (2022a, b). A remarkable consequence of this difference in symmetry is that the AM leads to momentum-dependent spin-splitting electronic band structures but maintains symmetry-compensated zero net magnetization. Thus, it is not only distinct from ferromagnetism which leads to momentum-independent spin splitting and finite magnetization, but also distinguishes itself from antiferromagnetism which results in degenerate band structures. Notably, the spin splitting induced by AM can reach the order of $1$eV, and the symmetry pattern of the spin splitting is rich, exemplified by the classification of AM into groups dubbed $d$-, $g$- and $i$-wave AM. Excitingly, several materials, including insulating MnTeOsumi et al. (2024); Lee et al. (2024); Krempaský et al. (2024); Hajlaoui et al. (2024) and metallic CrSbReimers et al. (2024); Ding et al. (2024); Yang et al. (2024); Zeng et al. (2024); Li et al. (2024), have been experimentally confirmed to be altermagnets by high-resolution angle-resolved photoemission spectroscopy (ARPES). On the theoretical front, first-principle Guo et al. (2023); Gao et al. (2023); Qu et al. (2024); Sødequist and Olsen (2024); Bernardini et al. (2024); Sheoran and Bhattacharya (2024); Milivojević et al. (2024); Zeng and Zhao (2024); Liu et al. (2024) and model Leeb et al. (2024); Das et al. (2023); Sato et al. (2023); Roig et al. (2024); Cônsoli and Vojta (2024); Yu et al. (2024); Bose et al. (2024) calculations have predicted an ever growing list of candidates for AM. Furthermore, many studies have also shown that AM can give rise to numerous interesting effects and phases, such as giant and tunneling magnetoresistanceŠmejkal et al. (2022c), diverse tunneling phenomena in superconductor/altermagnet junctionsOuassou et al. (2023); Sun et al. (2023); Papaj (2023); Beenakker and Vakhtel (2023); Cheng and Sun (2024); Nagae et al. (2024); Gløckner Giil et al. (2024); Das and Soori (2024); Lu et al. (2024), finite-momentum Cooper pairingZhang et al. (2024); Sumita et al. (2023); Chakraborty and Black-Schaffer (2023), unconventional superconductivityZhu et al. (2023); Wei et al. (2024); Brekke et al. (2023); Mæland et al. (2024); Chourasia et al. (2024), various types of Hall effectsParshukov et al. (2024); Fang et al. (2023); Attias et al. (2024); Hoyer et al. (2024), and anisotropic RKKY interaction between spin impuritiesLee (2023); Amundsen et al. (2024).
Because altermagnetic materials are often grown on substrates which can induce spin-orbit coupling (SOC), theoretical studies of altermagnetic materials need to also take into account this effect. SOC is known as another basic mechanism giving rise to momentum-dependent spin splitting and the interplay of AM and SOC can result in band structures of rich topological propertiesFernandes et al. (2024). Thus far, the $d$-wave AM has been the focus of most of the theoretical studies. It has been shown that the band structure of an intrinsic $d$-wave altermagnet normally has spin-polarized Dirac points in two dimensions (2D), and the further presence of SOC can gap out the Dirac points and result in first-order topological insulator phasesAntonenko et al. (2024); Parshukov et al. (2024). Furthermore, when intrinsic superconductivity occurs in a 2D $d$-wave altermagnet with Rashba SOC, both first-order and second-order topological superconductivity are found to emergeZhu et al. (2023). Lastly, it has been shown that second-order topological insulators or superconductors can be obtained in hybrid systems composed of $d$-wave altermagnets and first-order topological insulators or superconductorsGhorashi et al. (2023); Li and Liu (2023); Li et al. (2024); Li (2024); Ezawa (2024). In all these studies, the nontrivial momentum-space topology of the bulk bands is manifested in the presence of topological boundary states, dictated by the bulk-boundary correspondence. Interestingly, real-space topological defects, which are ubiquitous in materials and are characterized by real-space topological invariants, can also reflect the momentum-space topology in a distinct way, known as the bulk-defect correspondenceTeo and Kane (2010); Chiu et al. (2016); Yao et al. (2017).
The types of topological defects in materials are diverseMermin (1979); Kleman and Friedel (2008). For lattice topological defects, disclinations and dislocations are two classes that attract particular interest in the context of topological phasesTeo and Hughes (2017); Lin et al. (2023); Asahi and Nagaosa (2012); Bühler et al. (2014); Hughes et al. (2014); Slager et al. (2014); de Juan et al. (2014); van Miert and Ortix (2018); Roy and Juričić (2021); Soto-Garrido et al. (2020); Schindler et al. (2022); Hu and Zhang (2024). In the seminal work of Ran, Zhang and VishwanathRan et al. (2009), it was discovered that a lattice dislocation can harbor a pair of 1D gapless helical modes in 3D topological insulators; the topological criterion for the existence of these topological dislocation modes is given by $\textbf{B}\cdot\textbf{M}_{\nu}=\pi\,(\text{mod}\,2\pi)$, where B refers to the Burgers vector which is the real-space topological invariant characterizing the dislocation, and $\textbf{M}_{\nu}=\sum_{i=1}^{3}\nu_{i}\textbf{G}_{i}/2$. Here $\nu_{i}=\{0,1\}$ is the weak topological indices defined on the high symmetry planes at the Brillouin zone’s boundaryFu and Kane (2007) and $\textbf{G}_{i}$ is the reciprocal lattice vectors. $\textbf{M}_{\nu}$ acts like a time-reversal invariant momentum. Later, it was demonstrated that the topological criterion can also be generalized to 2D and is applicable to topological superconductors as wellRan (2010). In 2D, when the topological criterion is fulfilled, the topological dislocation modes are 0D bound states. As an application, it was shown that the presence or absence of dislocation bound states in a 2D topological insulator with $C_{4z}$ rotation symmetry can serve as a bulk probe to diagnose the location of band inversionJuričić et al. (2012).
Figure 1: (a) A schematic diagram of field-sensitive dislocation bound states and their experimental measurement by scanning tunneling microscopy. The 2D altermagnetic material grown on an insulator substrate or an $s$-wave superconductor has two pairs of dislocations with perpendicular Burgers vectors. The color correspondence between the bell-shaped surfaces denoting topological dislocation bound states and the arrows denoting the exchange field’s direction illustrates the fact that which pair of the dislocations carries topological bound states relies on the exchange field’s direction. (b) Details of the dislocation configuration. The two blue (red) dots correspond to the cores of a pair of dislocations with Burgers vector $\textbf{B}=\pm\bm{e_{x}}$ (±$\pm\bm{e_{y}}$). The two pairs of dislocations with perpendicular Burgers vectors form a cross-shaped defect. (c) 2D Brillouin zone. $\boldsymbol{\Gamma}=(0,0)$, $\textbf{X}=(\pi,0)$, $\textbf{Y}=(0,\pi)$ and $\textbf{M}=(\pi,\pi)$ are four high symmetry points. Two weak topological indices, $\nu_{x}$ and $\nu_{y}$, which determine the presence or absence of topological dislocation bound states, are defined on the blue and red lines at the Brillouin zone’s boundary.
In this paper, we investigate topological dislocation modes in a 2D $d$-wave altermagnet grown on a substrate. Because of the structure asymmetry, Rashba SOC becomes an important factor in the determination of the band structure in the altermagnet. Although the $d$-wave AM breaks the time-reversal symmetry ($\mathcal{T}$) and $C_{4z}$ rotation symmetry, it respects their combination, the $C_{4z}\mathcal{T}$ symmetry. As the Rashba SOC also respects this symmetry, the cooperation of $d$-wave AM and Rashba SOC leads to a unique spin-splitting band structure with spin textures and Berry curvatures respecting this symmetry tooZhu et al. (2023). Furthermore, despite the existence of spin splitting at generic momenta in the Brillouin zone, the band structure has nodal points at the two $C_{4z}\mathcal{T}$ invariant momenta. The $C_{4z}\mathcal{T}$-symmetric band structure serves as the indication of a critical phase, since a weak perturbation breaking the $C_{4z}\mathcal{T}$ symmetry can gap out the nodal points and lead to topological gapped phases. As we are interested in these gapped phases, we consider the presence of an additional exchange field which will break the $C_{4z}\mathcal{T}$ symmetry. Remarkably, we find that both the strong topological index (Chern number) and the $Z_{2}$-valued weak topological indices characterizing the gapped band structure depend sensitively on the exchange field’s direction. As a result, whether a dislocation harbors topological bound states hinges on the exchange field’s direction. Furthermore, when the substrate is an $s$-wave superconductor, we find that a similar dependence of band topology on the exchange field gives rise to field-sensitive dislocation Majorana zero modes.
The rest of the paper is organized as follows. In Sec.II, we describe the theoretical model and analyze the dependence of topological indices on the exchange field. By considering a cross-shaped defect consisting of two pairs of dislocations with perpendicular Burgers vectors, we illustrate that the presence of topological bound states at the dislocation cores depends sensitively on the exchange field’s direction. In Sec.III, we generalize this analysis and show that field-sensitive dislocation Majorana zero modes can be obtained when the substrate is an $s$-wave superconductor. In Sec.IV, we discuss our findings and conclude the paper.
II Dislocation bound states in 2D metal with $d$-wave AM and Rashba SOC
We first consider a 2D altermagnet grown on an insulator substrate and subject to a perpendicular exchange/Zeeman field, as illustrated in Fig.1(a). In this paper, we do not distinguish between Zeeman field and exchange field in terms of terminology. In the absence of defects, the tight-binding Hamiltonian is given byZhu et al. (2023) $H=\sum_{\bm{k}}c_{\bm{k}}^{{\dagger}}\mathcal{H}_{0}(\bm{k})c_{\bm{k}}$ with $c_{\bm{k}}^{{\dagger}}=(c_{\bm{k},\uparrow}^{{\dagger}},c_{\bm{k},\downarrow}^{{\dagger}})$ and
$\mathcal{H}_{0}(\bm{k})$ $=$ $-2t(\cos{k_{x}}+\cos{k_{y}})\sigma_{0}+2\lambda(\sin{k_{y}}\sigma_{x}-\sin{k_{x}}\sigma_{y})$ (1)
$+[2t_{\rm AM}(\cos{k_{x}}-\cos{k_{y}})+M_{z}]\sigma_{z},$
Here $\sigma_{i}$ are Pauli matrices acting on the spin degrees of freedom; the first term in $\mathcal{H}_{0}$ is the kinetic energy arising from the nearest-neighbor hoppings, the second term denotes the Rashba SOC originating from structure asymmetry, and the last term accounts for the presence of two types of exchange fields, where the first momentum-dependent part ($t_{\rm AM}$) is attributed to $d$-wave AM and the second momentum-independent part ($M_{z}$) to an external perpendicular magnetic field or a ferromagnetic insulator substrate whose magnetization is perpendicular to the plane. Throughout the lattice constant is set to unity for notational simplicity, and without loss of generality the parameters $t$, $\lambda$ and $t_{\rm AM}$ are assumed to be non-negative for the convenience of discussion.
Figure 2: Energy spectrum for a sample of ribbon geometry. For the top row, the system takes periodic (open) boundary conditions in the $x$ ($y$) direction. The bottom row is just the opposite. In (a) and (b), $M_{z}=0.24$, $C_{-}=-1$. The chiral edge state propagates clockwise on the boundary, and the edge-state spectrum crossings occur at $k_{x}=\pi$ and $k_{y}=0$, indicating $(\nu_{x},\nu_{y})=(1,0)$. In (c) and (d), $M_{z}=-0.24$, $C_{-}=1$. The chiral edge state becomes anticlockwise, and the edge-state spectrum crossings are switched to $k_{x}=0$ and $k_{y}=\pi$, indicating $(\nu_{x},\nu_{y})=(0,1)$. The red and blue solid lines in (a) and (c) [(b) and (d)] refer to the chiral edge states on the top and bottom $y$-normal (left and right $x$-normal) edges. Common parameters are $t=0$, $t_{\rm AM}=0.3$ and $\lambda=0.3$.
When $M_{z}=0$, the Hamiltonian has the $C_{4z}\mathcal{T}$ symmetry, even though the $C_{4z}$ rotation symmetry ($C_{4z}=e^{i\frac{\pi}{4}\sigma_{z}}$) and time-reversal symmetry ($\mathcal{T}=-i\sigma_{y}\mathcal{K}$ with $\mathcal{K}$ the complex conjugate operator) are independently broken by the $d$-wave AM. Because of this combined symmetry, the two energy bands have Kramers degeneracies at the two $C_{4z}\mathcal{T}$-invariant momenta, i.e., $\bm{\Gamma}$ and MZhu et al. (2023). Once $M_{z}$ becomes finite, the $C_{4z}\mathcal{T}$ symmetry is broken and the band degeneracies at $\bm{\Gamma}$ and M are lifted, resulting in a finite energy gap between the two bands. As the system does not have time-reversal symmetry, the gapped band structure is characterized by the first-class Chern number. The Chern numbers characterizing the two bands are given byQi et al. (2006)
$C_{\pm}$ $=$ ±$\pm\frac{1}{2\pi}\int_{BZ}\frac{\bm{d}(\bm{k})\cdot[\partial_{k_{x}}\bm{d}(\bm{k})\times\partial_{k_{y}}\bm{d}(\bm{k})]}{2|\bm{d}(\bm{k})|^{3}}d^{2}k$ (2)
$=$ ±$\pm\begin{cases}{\rm sgn}(M_{z}),&0<|M_{z}|<4t_{\rm AM},\\ 0,&|M_{z}|>4t_{\rm AM}.\end{cases}$
where the subscript +/+/-+ / - refers to the upper/lower band, and $\bm{d}(\boldsymbol{k})=(2\lambda\sin k_{y},-2\lambda\sin k_{x},2t_{\rm AM}(\cos{k_{x}}-\cos{k_{y}})+M_{z})$, with the components $d_{i}(\boldsymbol{k})$ corresponding to the coefficient functions in front of the Pauli matrix $\sigma_{i}$ in Eq. (1). The dependence of Chern number on $M_{z}$ suggests that an arbitrarily weak perpendicular exchange field render the band structure topologically nontrivial. Furthermore, the Chern number has a sensitive dependence on the direction of the exchange field. Thus, a reversal of the direction of the exchange field will change the sign of the Chern number. As topological phases have bulk-boundary correspondence, it is natural to expect that the topological boundary states will also have an intriguing dependence on the exchange field.
By calculating the energy spectrum for a sample of ribbon geometry, we find as expected that the chiral edge state reverses its chirality when the exchange field reverses its direction. What is surprising, however, is that the momentum at which the edge-state spectrum crosses undergoes a jump. When $M_{z}>0$, the edge-state spectrum crossing occurs at $k_{x}=\pi$ on the $y$-normal edges [Fig.2(a)] and at $k_{y}=0$ on the $x$-normal edges [Fig.2(b)]. Reversing the direction of the exchange field, the crossing momentum has a jump of half the reciprocal lattice vector. That is, the edge-state spectrum crossing is shifted to $k_{x}=0$ on the $y$-normal edges [Fig.2(c)] and to $k_{y}=\pi$ on the $x$-normal edges [Fig.2(d)]. The edge-state spectrum crossing is connected to weak topological indices defined on one-dimensional lower submanifolds of the Brillouin zone, namely, the high symmetry lines of the Brillouin zone. The jump of the crossing momentum then suggests that the weak topological indices of this Hamiltonian also have a sensitive dependence on the direction of the exchange field.
The weak topological indices of interest in this work are defined on the two high symmetry lines at the boundaries of the 2D Brillouin zone, as illustrated in Fig.1(c). The two weak topological indices, labeled as $\nu_{x}$ and $\nu_{y}$, are $Z_{2}$-valued and their parities determine whether there is an odd number of edge-state spectrum crossings at $k_{x}=\pi$ and $k_{y}=\pi$ of the respective boundary Brillouin zone. Although the $C_{4z}\mathcal{T}$ symmetry is broken by the exchange field, the whole system still has the $C_{2z}$ rotation symmetry described by the symmetry operator $C_{2z}=i\sigma_{z}$. Owing to the existence of this crystalline symmetry, the weak topological indices can be defined in terms of the eigenvalues of the $C_{2z}$ operator at $C_{2z}$-invariant momenta. Following the same spirit as the Fu-Kane formula for $Z_{2}$ invariants of topological insulatorsFu and Kane (2007), the explicit formulas for the two $Z_{2}$-valued weak topological indices are given by
$(-1)^{\nu_{x}}=-\xi(\textbf{X})\xi(\textbf{M}),(-1)^{\nu_{y}}=-\xi(\textbf{Y})\xi(\textbf{M}).$ (3)
where $\xi(\bm{k}_{R})$ is the $C_{2z}$ operator’s eigenvalue for the lower-band eigenstate at the $C_{2z}$-invariant momentum $\bm{k}_{R}\in\{\textbf{X},\textbf{Y},\textbf{M}\}$, i.e., $C_{2z}|u(\bm{k}_{R})\rangle=\xi(\bm{k}_{R})|u(\bm{k}_{R})\rangle$. Because these eigenvalues take values of ±$\pm i$, a factor of “$-1$” is introduced on the right hand side of the above equations. A straightforward calculation reveals
$(\nu_{x},\nu_{y})=\left\{\begin{array}[]{cc}(1,0),&0<M_{z}<4t_{\rm AM},\\ (0,1),&-4t_{\rm AM}<M_{z}<0,\\ (0,0),&|M_{z}|>4t_{\rm AM}.\end{array}\right.$ (7)
The above result shows explicitly that the two weak topological indices switch their values when the exchange field in the weak-field regime ($|M_{z}|<4t_{\rm AM}$) reverses its direction. Based on the two weak topological indices, the momentum $\textbf{M}_{\nu}$ reflecting the band topology is given by $\textbf{M}_{\nu}=\pi(\nu_{x}\bm{e_{x}}+\nu_{y}\bm{e_{y}})$.
Figure 3: Phase diagram determined via weak topological indices. The red and blue regions correspond to $(\nu_{x},\nu_{y})=(0,1)$ and $(1,0)$, respectively. The two insets below the phase diagram are energy spectra for a system with periodic boundary conditions in both $x$ and $y$ directions. The results show that, in the region with $(\nu_{x},\nu_{y})=(0,1)$, dislocations with Burgers vector $\textbf{B}=\pm\bm{e_{y}}$ harbor bound states, while the ones with $\textbf{B}=\pm\bm{e_{x}}$ do not, and the situation is just the opposite for the region with $(\nu_{x},\nu_{y})=(1,0)$. For the left (right) inset, $t_{\rm AM}=0.3$ and $M_{z}=-0.24$ $(0.24)$. The values of other parameters are $t=0$ and $\lambda=0.3$.
Let us now consider the presence of dislocations in the system. As aforementioned, the topological criterion for the presence of topological bound states at dislocations in 2D is also $\textbf{B}\cdot\textbf{M}_{\nu}=\pi$ (mod $2\pi$)Ran (2010). Because of the $\textbf{M}_{\nu}$’s sensitive dependence on the exchange field’s direction, whether a dislocation with a fixed Burgers vector carries a topological bound state will thereby also sensitively depend on the exchange field’s direction.
To verify this expectation, we place a cross-shaped defect consisting of two pairs of dislocations in the system [see Fig.1(b)] and numerically diagonalize the Hamiltonian under periodic boundary conditions in both $x$ and $y$ directions. In the case of $0<M_{z}<4t_{\rm AM}$ for which $(\nu_{x},\nu_{y})=(1,0)$, we find that the pair of dislocations with Burgers vector $\textbf{B}=\pm\bm{e_{x}}$ harbor topological bound states while those with Burgers vector $\textbf{B}=\pm\bm{e_{y}}$ do not, as shown in Fig.3. The picture is just reversed when $-4t_{\rm AM}<M_{z}<0$, demonstrating that the presence or the absence of topological bound states at a dislocation can be tuned by simply adjusting the exchange field’s direction.
III Field-sensitive dislocation Majorana zero modes
Controllability is a desired property in the application of Majorana zero modes in topological quantum computationAasen et al. (2016). Following the same line of argument as before, we will show in this section that field-sensitive Majorana zero modes can also be achieved by taking advantage of the AM. To be specific, we now consider the senario where a 2D $d$-wave altermagnetic metal is grown on a fully-gapped $s$-wave superconductor [see Fig.1(a)] and is assumed to inherit the $s$-wave superconductivity from the bulk superconductor though the proximity effect. Within the Bogoliubov-de Gennes (BdG) framework, the effective Hamiltonian is given by $H=\frac{1}{2}\sum_{\bm{k}}\Psi_{\bm{k}}^{\dagger}\mathcal{H}_{\rm BdG}(\bm{k})\Psi_{\bm{k}}$, where $\Psi^{\dagger}(\bm{k})=(c^{\dagger}_{\bm{k},\uparrow},c^{\dagger}_{\bm{k},\downarrow},c_{-\bm{k},\uparrow},c_{-\bm{k},\downarrow})$ and
$\mathcal{H}_{\rm BdG}(\bm{k})$ $=$ $[-2t(\cos{k_{x}}+\cos{k_{y}})-\mu]\tau_{z}\sigma_{0}$ (8)
$+[2t_{\rm AM}(\cos{k_{x}}-\cos{k_{y}})+M_{z}]\tau_{z}\sigma_{z}$
$+2\lambda(\sin{k_{y}}\tau_{0}\sigma_{x}-\sin{k_{x}}\tau_{z}\sigma_{y})+\Delta_{s}\tau_{y}\sigma_{y}.$
Here $\mu$ is the chemical potential, $\Delta_{s}$ is the proximity-induced $s$-wave pairing amplitude and the new set of Pauli matrices $\tau_{i}$ acts on the particle-hole degrees of freedom.
We again first focus on the band topology of the BdG Hamiltonian. Since the $s$-wave pairing does not break the time-reversal symmetry and $C_{4z}$ rotation symmetry, the BdG Hamiltonian also respects the $C_{4z}\mathcal{T}$ symmetry when $M_{z}=0$. In a previous work, we have shown that the $C_{4z}\mathcal{T}$ symmetry forbids a 2D gapped superconductor to have a nonzero Chern numberZhu et al. (2023). Despite the absence of strong topology characterized by the Chern number, weak topology is compatible with this symmetry and can be nontrivial when the system’s parameters fulfill certain conditionsGhorashi et al. (2023). To see this, we again make use of the eigenvalues of the $C_{2z}$ operator, which is now given by $C_{2z}=-i\tau_{z}\sigma_{z}$ due to the inclusion of particle-hole degrees of freedom. More specifically, we define
$(-1)^{\nu_{x}}$ $=$ $(-1)^{N}\prod_{n=1}^{N}\xi_{n}(\textbf{X})\xi_{n}(\textbf{M}),$
$(-1)^{\nu_{y}}$ $=$ $(-1)^{N}\prod_{n=1}^{N}\xi_{n}(\textbf{Y})\xi_{n}(\textbf{M}),$ (9)
where $N=2$ denotes the number of bands below $E=0$, $n$ is the band index with $E_{n}<0$, and $\xi_{n}(\bm{k}_{R})$ is the $C_{2z}$ operator’s eigenvalue for the negative-energy eigenstate at $\bm{k}_{R}\in\{\textbf{X},\textbf{Y},\textbf{M}\}$, i.e., $C_{2z}|u_{n}(\bm{k}_{R})\rangle=\xi_{n}(\bm{k}_{R})|u_{n}(\bm{k}_{R})\rangle$. A straightforward calculation yields the weak topological indices at $M_{z}=0$,
$(\nu_{x},\nu_{y})=\left\{\begin{array}[]{cc}(1,1),&4t_{\rm AM}>\sqrt{\mu^{2}+\Delta_{s}^{2}},\\ (0,0),&4t_{\rm AM}<\sqrt{\mu^{2}+\Delta_{s}^{2}}.\end{array}\right.$ (12)
This result indicates that, despite the impossibility of chiral Majorana modes when $M_{z}=0$, Majorana zero modes can be created at dislocations with Burgers vectors equal to either ±$\pm\bm{e_{x}}$ or ±$\pm\bm{e_{y}}$. Before explicitly showing this, we first complete the analysis of the bulk topology when $M_{z}\neq 0$.
Figure 4: BdG energy spectrum for a sample of ribbon geometry. The system takes open boundary directions in the $y$ ($x$) direction and periodic boundary conditions in the $x$ ($y$) direction for the top (bottom) row. In (a) and (b), $M_{z}=0$, $C=0$, and the locations of edge-state spectrum crossings suggest $(\nu_{x},\nu_{y})=(1,1)$. In (c) and (d), $M_{z}=-0.3$, $C=-1$, and the locations of edge-state spectrum crossings suggest $(\nu_{x},\nu_{y})=(1,0)$. In (e) and (f), $M_{z}=0.3$, $C=1$, and the locations of edge-state spectrum crossings suggest $(\nu_{x},\nu_{y})=(0,1)$. Common parameters are $t=0.2$, $t_{\rm AM}=0.3$, $\lambda=0.3$, $\Delta_{s}=0.2$, and $\mu=-1.1$.
Similar to the previous nonsuperconducting case, a finite $M_{z}$ breaks the $C_{4z}\mathcal{T}$ symmetry and can induce topological superconducting phases characterized by nonzero Chern numbers. The topological phase diagram can be easily determined since the change of Chern number is associated with the close of bulk energy gapSau et al. (2010), which takes place when $M_{z}=\{\pm\sqrt{(4t\pm\mu)^{2}+\Delta_{s}^{2}}$, ±$\pm 4t_{\rm AM}\pm\sqrt{\mu^{2}+\Delta_{s}^{2}}$}. Here we are not interested in determining the complete phase diagram. Rather, we are interested in the region where the exchange field is weak, i.e., when $M_{z}$ is comparable to $\Delta_{s}$, and when the chemical potential is close to the critical points between the weak topological superconductor and trivial superconductors, i.e., when $\mu$ is close to $\mu_{c,\pm}=\pm\sqrt{16t_{\rm AM}^{2}-\Delta_{s}^{2}}$. Without loss of generality, we take $\mu=\mu_{c,-}+\delta$ where $\delta$ is a small real positive constant. Then, increasing the exchange field’s strength from zero, the energy gap closes at X when $M_{z}=4t_{\rm AM}-\sqrt{\mu^{2}+\Delta_{s}^{2}}\simeq\delta$ (assuming $\Delta_{s}\ll t_{\rm AM}$), and at Y when $M_{z}=-4t_{\rm AM}+\sqrt{\mu^{2}+\Delta_{s}^{2}}\simeq-\delta$. The close of energy gap at X changes the total Chern number of the bands below $E=0$ from $C=0$ to $C=1$, and the weak topological indices $(\nu_{x},\nu_{y})$ from $(1,1)$ to $(0,1)$. In contrast, the close of energy gap at Y changes the total Chern number of the bands below $E=0$ from $C=0$ to $C=-1$, and the weak topological indices $(\nu_{x},\nu_{y})$ from $(1,1)$ to $(1,0)$.
In Fig.4, we show the energy spectrum for a sample of ribbon geometry. In Figs.4(a) and 4(b), we see that there are two counter-propagating gapless states on one edge and the edge-state spectrum crossing occurs at time-reversal invariant momenta on both $x$-normal and $y$-normal edges; this is consistent with $C=0$ and $(\nu_{x},\nu_{y})=(1,1)$ for $M_{z}=0$ and $4t_{\rm AM}>\sqrt{\mu^{2}+\Delta_{s}^{2}}$. In Figs.4(c) and 4(d), the number of chiral edge states and the locations of edge-state spectrum crossings suggest that $C=-1$ and $(\nu_{x},\nu_{y})=(1,0)$, which is also consistent with the values as analyzed above. In Figs.4(e) and 4(f), the results show that a reversal of the exchange field’s direction reverses the chirality of the gapless edge state, and leads to a switching of the locations of edge-state spectrum crossings on the $x$-normal and $y$-normal edges. This again agrees with the sign change of the Chern number and the switching of the values of $\nu_{x}$ and $\nu_{y}$ accompanying this process.
Placing a cross-shaped defect in the superconducting system and diagonalizing the Hamiltonian under periodic boundary conditions in both directions, we find the presence of Majorana zero modes at dislocations when the topological criterion $\textbf{B}\cdot\textbf{G}_{\nu}=\pi\,(\text{mod}\,2\pi)$ is fulfilled. To be specific, when the system is a weak topological superconductor with $(\nu_{x},\nu_{y})=(1,1)$ (indicated by the purple region of Fig.5), we find that there are four Majorana zero modes with their wave functions localized at the cores of the dislocations, as shown in the inset on top of the purple region. The result suggests that dislocations with Burgers vector $\textbf{B}=\pm\bm{e_{x}}$ and ±$\pm\bm{e_{y}}$ all harbor Majorana zero modes at their cores, which agrees with the topological criterion. In comparison, when $\nu_{x}$ ($\nu_{y}$) is trivialized by the exchange field, the two Majorana zero modes at the dislocations with $\textbf{B}=\pm\bm{e_{x}}$ (±$\pm\bm{e_{y}}$) disappear, while the two Majorana zero modes at the dislocations with $\textbf{B}=\pm\bm{e_{y}}$ (±$\pm\bm{e_{x}}$) remain intact; this is clearly shown in the inset on top of the red (blue) region in Fig.5. We can draw an important conclusion from Fig.5. Namely, when the chemical potential is close to $\mu_{c,-}$, tuning the direction of a weak exchange field can control the presence or absence of Majorana zero modes at a dislocation. These results suggest that the unique spin-splitting band structure induced by AM and Rashba SOC provides a basis for the realization of controllable Majorana modes.
Figure 5: Phase diagram around the critical point with $\mu_{c,-}=-\sqrt{16t_{\rm AM}^{2}-\Delta_{s}^{2}}$ and $M_{c}=0$. The phase boundaries are given by $M_{z}=\pm(4t_{\rm AM}-\sqrt{\mu^{2}+\Delta_{s}^{2}})$. The weak topological indices $(\nu_{x},\nu_{y})$ in the blue, white, red and purple regions take values of $(1,0)$, $(0,0)$, $(0,1)$ and (1,1), respectively. The four insets at the corresponding regions show the number and distributions of Majorana zero modes at the two pairs of dislocations with $\textbf{B}=\pm\bm{e_{x}}$ and ±$\pm\bm{e_{y}}$. The corresponding parameters of the insets on the left, right, top and bottom are $(\mu,M_{z})=(-1.183,-0.3)$, $(\mu,M_{z})=(-1.183,0.3)$, $(\mu,M_{z})=(-0.6,0)$ and $(\mu,M_{z})=(-1.4,0)$. Common parameters are $t=0.2$, $t_{\rm AM}=0.3$, $\lambda=0.3$ and $\Delta_{s}=0.2$.
IV Discussions and conclusions
The combination of $d$-wave AM and Rashba SOC leads to a unique spin-splitting band structure respecting the $C_{4z}\mathcal{T}$ symmetry. Breaking the $C_{4z}\mathcal{T}$ symmetry by an exchange field, we find that both the strong topology and the weak topology of the resulting band structure show a sensitive dependence on the exchange field’s direction. As a consequence of the sensitive dependence of weak topological indices on the exchange field’s direction, we find that the presence or absence of topological bound states at a dislocation can be easily controlled by adjusting the exchange field’s direction. By putting the 2D altermagnetic metal in proximity to an $s$-wave superconductor, we find a similar sensitive dependence of the band topology and dislocation bound states on the exchange field when the chemical potential is appropriately chosen. As the dislocation bound states are Majorana zero modes in this case, their sensitivity towards external fields implies high degrees of controllability; this may make the superconducting altermagnetic materials stand out as platforms to detect and manipulate Majorana zero modes.
From a symmetry perspective, the change from one set of topological dislocation bound states into their $C_{4z}$-rotational counterparts upon reversing the exchange field’s direction is a direct manifestation of the $C_{4z}\mathcal{T}$ symmetry of the pristine altermagnetic system. In the light of this, an observation of this switching behavior of topological dislocation bound states can be taken as a strong signature of $d$-wave AM. In experiments, the field-sensitive dislocation bound states are robust due to their topological origin and can be easily detected by scanning tunneling microscopy. Given the ubiquitous presense of topological defects in real materials and a continuously accumulated list of candidates for altermagnetic materials, our findings unveil a promising route to an effective experimental diagnosis of AM in these materials.
Acknowledgements
D. Z., D. L., Z.-Y. Z, and Z. Y. are supported by the National Natural Science Foundation of China (Grant No. 12174455), Natural Science Foundation of Guangdong Province (Grant No. 2021B1515020026), and Guangdong Basic and Applied Basic Research Foundation (Grant No. 2023B1515040023). Z. W. is supported by National Key R&\&&D Program of China (Grant No. 2022YFA1404103), NSFC (Grant No. 11974161) and Shenzhen Science and Technology Program (Grant No. KQTD20200820113010023).